# literature.bib

### v2

@@ Abstract
We analyze a recently proposed algorithm for the problem of sampling from probability distributions $\mu^{\ast}$ in $\mathbb{R}^{d}$ with a Lebesgue density of the form $\mu^{\ast}(x)\propto\exp(-f(Kx)-g(x))$, where $K$ is a linear operator and $f,g$ convex and non-smooth. The algorithm is a generalization of the primal-dual hybrid gradient (PDHG) convex optimization algorithm to a sampling scheme. We analyze the method’s continuous time limit, an SDE in the joint primal-dual variable. We give mild conditions under which the corresponding Fokker-Planck equation converges to a unique stationary state, which however does not concentrate in the dual variable and consequently does not have $\mu^{\ast}$ as its primal marginal. Under a smoothness assumption on $f$, we show that the scheme converges to the purely primal overdamped Langevin diffusion in the limit of small primal and large dual step sizes. We further prove that the target can never be the primal marginal of the invariant solution for any modification of the SDE with space-homogeneous diffusion coefficient. A correction with inhomogeneous diffusion coefficient and the correct invariant solution is proposed, but the scheme requires the same smoothness assumptions on $f$ and is numerically inferior to overdamped Langevin diffusion. We demonstrate our findings numerically first on small-scale examples in which we can exactly verify the theoretical results, and subsequently on typical examples of larger scale from Bayesian imaging inverse problems.
@@ 1 Introduction
We analyze a recently proposed algorithm for sampling from a probability distribution on $\mathcal{X}\coloneqq\mathbb{R}^{d}$ with log-concave, non-smooth density. In particular, we are interested in sampling from target distributions with Lebesgue densities of the form
@@
A standard class of algorithms to draw samples in Bayesian inference are Markov chain Monte Carlo (MCMC) methods. Compared to other types of MCMC methods, sampling schemes based on Langevin diffusion processes have proved efficient in inverse problems and imaging applications due to their preferable scaling behaviour to high dimensions [Roberts1996, Durmus2022, Durmus2019]. MCMC sampling algorithms based on Langevin diffusion typically perform a time discretisation of the stochastic differential equation (SDE)
$\mathrm{X}_{t}=\nabla\log\mu^{\ast}(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},$ (1.2)
@@
this iteration is also known as unadjusted Langevin algorithm (ULA). ULA and also similar implicit or semi-implicit time discretizations introduce a bias between the law of the samples and the target measure, hence many works have been concerned with the characterization of the invariant distribution of the Markov chain $(X^{n})$ and its bias [Roberts1996, Durmus2017, Wibisono2018, Dalalyan2019]. Further questions typically concern preconditioning and convergence speed of the Markov chain and the behaviour of Metropolis-Hastings correction steps to overcome the bias and draw unbiased samples from $\mu^{\ast}$.
In this work, we are concerned with a new type of Langevin sampling algorithm which has been proposed as Unadjusted Langevin Primal-Dual Algorithm (ULPDA) in [Narnhofer2022] and has since been mentioned in [Lau2023, Habring2023]. Unlike ULA or its variants, ULPDA can not be interpreted as a time discretization of (1.2). The algorithm consists in carrying out steps of a primal-dual optimization method, where the same type of stochasticity as in unadjusted Langevin algorithms is added to the primal variable in each step. Its motivation is hence inspired by the optimization view: Just as (1.3) can be interpreted as a gradient descent with an additive stochastic component, one might be able to obtain more general sampling algorithms by adding the right amount of stochasticity to more general classes of first-order optimization methods. This is, of course, not a rigorous argument for convergence, and until now, to the best of the authors’ knowledge, no convergence theory for ULPDA has been provided. Indeed, the situation seems to be more intricate: When we run ULPDA on a simple toy example, the distribution of the samples seems to depend more critically on the choice of step size hyperparameters $\tau,\sigma$ of the algorithm than it is known from ULA, see Figure 1.
@@ 1.1 Related Works
There is a vast literature on unadjusted Langevin algorithms, where we use the term to refer not only to (1.3), but any time discretization of (1.2). The analysis is based on the fact that the target distribution is the stationary distribution of the SDE (1.2) and that the dynamics converge to the target in continuous time [Roberts1996]. This is based on the fact that the dynamics can be understood as a gradient flow of relative entropy with respect to the target in Wasserstein space [Jordan1998, Wibisono2018]. Standard convergence results on ULA typically require a Lipschitz continuous gradient of the potential in order to bound the error introduced by time discretization [Dalalyan2019, Durmus2019, Durmus2019a].
Since non-smooth densities do not satisfy this gradient Lipschitz condition, many works have treated the extension to the non-smooth case. Generally, the approaches can be divided into two classes. The first regularizes non-smooth terms, resulting in a modified, smooth target and applies typical ULA to this modified distribution. The challenge then typically consists in bounding the terms between the target and the modified distribution, and numerically in trading off the introduced bias and convergence speed via the smoothing parameter [Pereyra2016, Durmus2022, Durmus2018, Lau2022, Cai2023]. The second approach consists in finding different discretizations of (1.2), i.e. implicit or semi-implicit schemes, that are implementable and provably convergent. These schemes are usually referred to as proximal or implicit Langevin methods, see e.g. [Pereyra2016, Bernton2018, Salim2020, Klatzer2023, Ehrhardt2023]. A work of particular relevance for our setting is [Habring2023], since the authors assume the same type of splitting of the potential $h=f\circ K+g$ as we do. Their method ‘Prox-Sub’ can be understood as a first-order forward-backward discretization of (1.2), with the forward step in $f\circ K$ realized as a subgradient since $f$ is assumed non-smooth. We will see that this method is intimately connected to the primal-dual sampling algorithm since it forms the limiting case in which the primal-dual stationary distribution can be related to the target. In [Narnhofer2022], the authors proposed the primal-dual scheme that we consider here, in order to generalize from the class of methods that employ purely primal discretizations of (1.2). A partial dualization of the potential is also considered in [Cai2023], although their analysis still requires the relevant part $f$ of the potential to be smooth. We also refer to [Lau2023] for a recent comparison of current methods applicable in the non-smooth case.
@@ 1.2 Contributions
  • For the continuous time limit of the iteration, we prove a contraction result using a coupling argument, which makes the dynamics asymptotically stable and guarantees convergence to the corresponding stationary solution.
  • We show that the marginal with respect to the primal variable of the stationary solution in continuous time is not the target distribution. This can be interpreted as being a consequence of the dynamics not concentrating in the dual variable, although the diffusion coefficient is degenerate with the stochastic term only added to the primal variable.
@@
  • We consider two potential frameworks for “correcting” the bias between continuous time stationary and target distribution. A proposed modification of the dynamics with non-homogeneous diffusion coefficient can be employed to force the target to become the stationary solution. Secondly, we consider the unmodified scheme, but with the ratio of primal and dual step sizes in the algorithm growing to infinity. In the limit, the stationary distribution of the unmodified scheme is shown to converge to the target under regularity assumptions on the dualized potential.
@@
  • In discrete time, we show stability of the algorithm in the sense that Wasserstein distances between two different instances are bounded if the potentials are convex. Under stronger regularity assumptions on $f^{\ast},g$, there exists a stationary solution to which the distribution of the iterates converges.
The manuscript is organized as follows. In Section 2, we give the considered sampling procedure based on primal-dual optimization and derive its continuous time limit SDE. Since this limiting SDE differs from (1.2), we devote Section 3 to the analysis of the continuous time setting. We derive the mean-field limit Fokker-Planck equation and show contraction properties in Section 3.1. In Section 3.2, we prove a central ‘negative’ result, i.e. that the target can not be the stationary distribution for any homogeneous diffusion coefficient and illustrate the (non-)concentration of the mean-field limit, which is crucial to understand how its stationary state relates to the target distribution. This relation is proven in Section 3.3, which allows to bound the bias due to the dualization in Wasserstein distance. We further show that the PDE can be modified to force the target to be its stationary distribution by employing a non-homogeneous diffusion coefficient. By transferring some of the time continuous arguments to the discretized algorithm and using typical computations known from primal-dual optimization methods, we show stability of the algorithm and convergence to a stationary state in Section 4. We present numerical illustrations of our most important findings in Section 5
2 The Primal-Dual Sampling Scheme
@@
2.1 Motivation for ULPDA
@@
The primal-dual optimization method (primal descent and dual ascent with extrapolation) for (2.2) takes the form
@@
$\left\{\begin{array}[]{l}Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n})\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})\end{array}\right.$ (2.3)
with step sizes $\tau$ and $\sigma$ and a relaxation parameter $\theta$ chosen accordingly. Herein, the proximal operator for a convex functional $\varphi$ and a step size $\gamma>0$ is defined by
$\operatorname{prox}_{\gamma\varphi}(z)\coloneqq(id+\gamma\partial\varphi)^{-1}(z)=\operatorname*{arg\,min}_{w}\left\{\varphi(w)+\frac{1}{2\gamma}\lVert w-z\rVert_{2}^{2}\right\}.$
The minimum is attained by a unique point, and the application of the proximal operator can be viewed as an implicit time step in the gradient flow of the functional $\varphi$. For convergence theory of the scheme (2.3) and possible accelerations we refer to [Chambolle2011, Chambolle2016].
In order to draw the parallels between computing the MAP of $\mu^{\ast}$ and drawing samples from $\mu^{\ast}$, we want to recall the standard unadjusted Langevin algorithm (1.3). Under appropriate conditions on $h$ and the step size $\tau$, ULA converges to a stationary distribution whose distance to the target $\mu^{\ast}$ can be bounded by a factor depending on $h$ and the considered metric on the space of probability distributions [Dalalyan2019]. ULA can be interpreted as a scheme alternating a gradient descent step in $h$ and a diffusion step (or, equivalently, a time step along the gradient flow of entropy in the Wasserstein space). In particular, the additional diffusion step ensures the samples have the correct covariance (up to a bias scaling linearly in the step size $\tau$).
@@
It appears natural to take a similar perspective for primal-dual optimization methods and add a diffusion step at some point in the iteration. Such a scheme was proposed (without further theoretical analysis) under the name unadjusted Langevin primal-dual algorithm (ULPDA) in [Narnhofer2022]. The iteration takes the form
@@
$\left\{\begin{array}[]{l}Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ \xi^{n+1}\sim\mathrm{N}(0,I_{d})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (2.4)
where the diffusion step is added in the primal variable $X_{k}$ and the overrelaxation parameter $\theta$ was set to $\theta=1$ in [Narnhofer2022] in the spirit of primal-dual optimization methods. Note that the stochastic term in the update for $X^{n+1}$ may be added in the argument of the proximal operator (as in [Narnhofer2022]) or after applying the proximal operator (as in 2.4). For the considerations in continuous time, this makes no difference, however it will simplify some of our analysis in discrete time based on couplings in Section 4.

### v3

@@ Abstract
We analyze a recently proposed class of algorithms for the problem of sampling from probability distributions $\mu^{\ast}$ in $\mathbb{R}^{d}$ with a Lebesgue density of the form $\mu^{\ast}(x)\propto\exp(-f(Kx)-g(x))$, where $K$ is a linear operator and $f,g$ convex and non-smooth. The method is a generalization of the primal-dual hybrid gradient optimization algorithm to a sampling scheme. We give the iteration’s continuous time limit, a stochastic differential equation in the joint primal-dual variable, and its mean field limit Fokker-Planck equation. Under mild conditions, the scheme converges to a unique stationary state in continuous and discrete time. Contrary to purely primal overdamped Langevin diffusion, the stationary state in continuous time does not have $\mu^{\ast}$ as its primal marginal. Thus, further analysis is carried out to bound the bias induced by the partial dualization, and potentially correct for it in the diffusion. Time discretizations of the diffusion lead to implementable algorithms, but, as is typical in Langevin Monte Carlo methods, introduce further bias. We prove bounds for these discretization errors, which allow to give convergence results relating the produced samples to the target. We demonstrate our findings numerically first on small-scale examples in which we can exactly verify the theoretical results, and subsequently on typical examples of larger scale from Bayesian imaging inverse problems.
Keywords: Langevin sampling, Markov chain Monte Carlo, Fokker-Planck equations, Coupling analysis, Primal-dual algorithms, Bayesian inference, Inverse problems
MSC codes: 35Q84, 47A52, 49N45, 60H10, 62F15, 65J22, 68U10
@@ 1 Introduction
We analyze a recently proposed class of algorithms for sampling from a probability distribution on $\mathcal{X}\coloneqq\mathbb{R}^{d}$ with log-concave, non-smooth density. We are interested in target distributions with Lebesgue densities of the form
@@
A standard class of algorithms to draw samples in Bayesian inference are Markov chain Monte Carlo (MCMC) methods. Compared to other types of MCMC methods, sampling schemes based on Langevin diffusion processes have proved efficient in inverse problems and imaging applications due to the preferable scaling behaviour to high dimensions [Roberts1996, Durmus2022, Durmus2019]. Typically, MCMC sampling algorithms based on Langevin diffusion perform a time discretisation of the stochastic differential equation (SDE)
$\mathrm{d}X_{t}=\nabla\log\mu^{\ast}(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t},$ (1.2)
@@
this iteration is also known as unadjusted Langevin algorithm (ULA). ULA and other time discretizations of (1.2) introduce a bias between the law of the samples and the target $\mu^{\ast}$, hence many works have been concerned with the characterization of the invariant distribution of the Markov chain $(X^{n})_{n}$ and its bias [Roberts1996, Durmus2017, Wibisono2018, Dalalyan2019]. Other typical questions concern preconditioning and convergence speed of the Markov chain and the behaviour of Metropolis-Hastings correction steps to overcome the bias and draw unbiased samples from $\mu^{\ast}$.
In this work, we are concerned with a new type of Langevin sampling algorithm which has been proposed as Unadjusted Langevin Primal-Dual Algorithm (ULPDA) in [Narnhofer2022] and has since been mentioned in [Lau2023, Habring2023]. Unlike ULA or its variants, ULPDA can not be interpreted as a time discretization of (1.2). It rather consists in steps of a primal-dual optimization method, where the same stochastic term $\sqrt{2\tau}\xi$ as in ULA is added to the primal iterate. The iteration takes the form
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I_{d})\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (1.4)
where $\tau,\sigma>0$ are step sizes, $\theta>0$ is an overrelaxation parameter, $f^{\ast}$ denotes the convex conjugate of $f$ and the proximal mapping of a convex function $\varphi$ is the unique value $\operatorname{prox}_{\gamma\varphi}(z)\coloneqq(id+\gamma\partial\varphi)^{-1}(z)$.
The motivation for this update comes from the optimization perspective: The ULA iteration (1.3) can be interpreted as gradient descent with an additive stochastic term. While ULA requires to know the gradients of the potentials, the advantage of primal-dual algorithms is that the iteration can be applied without the need to evaluate any gradients or even the proximal mapping of the composition $f\circ K$. $f$ and $g$ need only be convex, not necessarily differentiable. Hence, it appears desirable to obtain more general sampling algorithms by adding stochasticity to other first-order optimization methods like the primal-dual update.
This is, of course, not a rigorous argument for convergence, and until now, to the best of the authors’ knowledge, no convergence theory for ULPDA has been provided. Indeed, the situation seems to be more intricate: When we run ULPDA on a simple toy example, see Figure 1, the distribution of the samples seems to depend more critically on the choice of step size hyperparameters $\tau,\sigma$ of the algorithm than it is known from ULA.
@@ 1.1 Related Works
There is a vast literature on unadjusted Langevin algorithms, where we use the term to refer not only to (1.3), but any time discretization of (1.2). The convergence analysis is based on the simple fact that the target distribution is the stationary distribution of the SDE (1.2) and that the dynamics converge to the target in continuous time [Roberts1996]. As an underlying concept, this convergence can be explained by drawing the connection between the diffusion dynamics and, equivalently, the gradient flow of relative entropy with respect to the target in Wasserstein space [Jordan1998, Wibisono2018]. Standard convergence results on ULA typically require a Lipschitz continuous gradient of the potential in order to bound the error introduced by time discretization [Dalalyan2019, Durmus2019, Durmus2019a].
Since non-smooth densities do not satisfy this gradient Lipschitz condition, many works have treated the extension to the non-smooth case. Generally, the approaches can be divided into two classes. The first regularizes non-smooth terms, resulting in a modified, smooth target and applies typical ULA to this modified distribution. The challenge then typically consists in bounding the terms between the target and the modified distribution, and numerically in trading off the introduced bias and convergence speed via the smoothing parameter [Pereyra2016, Durmus2022, Durmus2018, Lau2022, Cai2023]. The second approach consists in finding different discretizations of (1.2), i.e. implicit or semi-implicit schemes, that are implementable and provably convergent. These schemes are usually referred to as proximal or implicit Langevin methods, see e.g. [Pereyra2016, Bernton2018, Salim2020, Klatzer2023, Ehrhardt2023]. A work of particular relevance for our setting is [Habring2023], since the authors assume the same type of splitting of the potential $h=f\circ K+g$ as we do. Their method ‘Prox-Sub’ can be understood as a first-order forward-backward discretization of (1.2), with the forward step in $f\circ K$ realized as a subgradient since $f$ is assumed non-smooth. We will see that this method is intimately connected to the primal-dual sampling algorithm since it forms the limiting case in which the primal-dual stationary distribution can be related to the target. In [Narnhofer2022], the authors proposed the primal-dual scheme that we consider here, in order to generalize from the class of methods that employ purely primal discretizations of (1.2). A partial dualization of the potential is also considered in [Cai2023], but the scheme differs from the one considered here in that it runs an inner iteration to solve the dual optimality condition. Further, the algorithm in [Cai2023] assumes smoothness of $g$ since the updates require evaluation of $\nabla g$. We also refer to [Lau2023] for a recent comparison of current methods applicable in the non-smooth case.
@@
We also mention that due to the dualization, the primal-dual Langevin sampling strategy can be interpreted as an alternating stochastic descent-ascent scheme. There is a more general interest in the analysis of Langevin descent-ascent schemes in the context of saddle point problems, e.g. for the recovery of Nash equilibria, see [Wang2024]. Our results give answers on the convergence in continuous and discretized time for the case where the saddle point functional is of a Lagrangian form obtained from rewriting a convex optimization problem as a saddle point problem.
@@ 1.2 Contributions
Our main contributions are the following (see also Figure 2 for an overview).
  • For the algorithm’s continuous time limit, we prove a contraction result using a coupling argument. This guarantees the dynamics are asymptotically stable and converge to a stationary distribution $\pi_{\lambda}(x,y)$, where $(x,y)\in\mathcal{X}\times\mathcal{Y}\eqqcolon\mathcal{Z}$ in the joint space of primal and dual variable. Crucially, the distribution depends on $\lambda>0$, a constant coupling primal and dual time scales.
  • We show that the marginal of $\pi_{\lambda}$ in the primal variable $x$ is not the target $\mu^{\ast}$. This can be interpreted as being a consequence of $\pi_{\lambda}$ not concentrating its mass on a lower-dimensional submanifold of $\mathcal{X}\times\mathcal{Y}$, although the diffusion coefficient is degenerate with diffusion only acting directly on $x$. In the case of smooth $f^{\ast},g$, we prove this property called hypoellipticity.
@@
  • We consider two potential frameworks for “correcting” the bias between the stationary distribution $\pi_{\lambda}$ and $\pi^{\ast}$ (a concentrated push-forward of the target $\mu^{\ast}$ to $\mathcal{Z}$). Firstly, a proposed modification of the dynamics with non-homogeneous diffusion coefficient can be employed to force $\pi^{\ast}$ to be the stationary solution. The target becomes the stationary distribution, but the discretization of the diffusion SDE is impractical. Secondly, we consider the unmodified scheme, but with the ratio $\lambda$ of dual and primal step sizes in the algorithm growing to infinity. The limit recovers the purely primal overdamped Langevin diffusion. Under regularity assumptions on the dualized potential $f$, we show convergence of $\pi_{\lambda}$ to $\pi^{\ast}$.
@@
  • In discrete time, we show stability of the algorithm in the sense that Wasserstein distances between two different instances are bounded if the potentials are convex. Under stronger regularity assumptions on $f^{\ast},g$, there exists a stationary solution $\pi_{\lambda,\tau}$ to which the distribution of the iterates converges. We then provide a bound of the bias, i.e. a bound on the Wasserstein distance between $\pi_{\lambda}$ and $\pi_{\lambda,\tau}$. This allows distance bounds between the generated samples and $\mu^{\ast}$.
Primal-dual ($0<\lambda<\infty$)Primal ($\lambda=\infty$) Continuous time Contraction to $\pi_{\lambda}$3.4}, Inconsistency {3.5,3.6}, Non-concentration {3.10}, Corrected SDE {3.13}. Stationary solution $\pi^{\ast}=(id,\nabla f\circ K)_{\#}\mu^{\ast}$ with the target $\mu^{\ast}$. Discretized time Stability {4.1}, Contraction to $\pi_{\lambda,\tau}$4.2}. ULA and variants from other discretizations of (1.2). {3.11}{4.8}other works,e.g. [Habring2023] for $h=f\circ K+g$
Figure 2: Overview of the results in this paper. The central theorems are indicated in curled brackets.
The manuscript is organized as follows. In Section 2, we give the considered sampling procedure based on primal-dual optimization and derive its continuous time limit SDE. Since the continuous time limit of the algorithm does not coincide with overdamped Langevin diffusion, we devote Section 3 to the analysis of the continuous time setting and its mean field limit Fokker-Planck equation. The stability and convergence analysis of the algorithm in discrete time is carried out in Section 4. We present numerical illustrations of our most important findings in Section 5 and conclude in Section 6.
2 Primal-Dual Langevin Sampling Schemes
@@
2.1 Motivation from Primal-Dual Optimization
@@
The primal-dual optimization method (primal descent and dual ascent with overrelaxation in the primal variable) for (2.2) takes the form
@@
$\left\{\begin{array}[]{l}Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n}\right)\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (2.3)
with step sizes $\tau,\sigma$, overrelaxation parameter $\theta$ and an initial guess $(X^{0},Y^{0})$. By convention, it is set $X_{\theta}^{0}=X^{0}$. For convergence theory of the scheme (2.3) and possible accelerations we refer to [Chambolle2011, Chambolle2016].
In order to draw the parallels between computing the MAP of $\mu^{\ast}$ and drawing samples from $\mu^{\ast}$, note that ULA (1.3) can be viewed as a scheme alternating a gradient descent step in $h=-\log\mu^{\ast}$ and a diffusion step. The diffusion is scaled to ensure the samples have the correct covariance up to a bias which is due to the discretization error and is controlled by the step size $\tau$. Similar to ULA, several first order optimization algorithms have been transformed into a sampling algorithm by inserting a scaled diffusion step at some point in the iteration [Lau2023]. The choice of optimization step is fairly flexible as long as the resulting method can be viewed as an alternating scheme combining a descent step (i.e., a discretized step along the gradient flow of a potential energy) with the diffusion (an exact step along the gradient flow of entropy) [Wibisono2018].
@@
It appears natural to take a similar perspective on primal-dual optimization methods and add a diffusion step at some point in the iteration. Such a scheme, in the form of (1.4), was proposed under the name unadjusted Langevin primal-dual algorithm (ULPDA) in [Narnhofer2022]. In this variant of the algorithm, the diffusion step is added only in the primal variable. We will consider both (1.4) as well as the more general scheme
@@
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I_{d+m})\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}\left(Y^{n}+\sigma KX_{\theta}^{n+1}\right)+B_{Y}\xi^{n+1}\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n})+B_{X}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n}),\end{array}\right.$ (2.4)
with general constant coefficients $B_{X}\in\mathbb{R}^{d\times(d+m)}$ and $B_{Y}\in\mathbb{R}^{m\times(d+m)}$ for the stochastic component. Note the increased dimension of the stochastic term compared to (1.4), which is a special case of (2.4) with $B_{X}=(\sqrt{2}I_{d},0)$ and $B_{Y}=0$.
As a further variant of the algorithm, we remark that in (1.4), we add the stochastic term $\sqrt{2\tau}\xi^{n+1}$ after applying the proximal operator. Similar to other proximal Langevin algorithms, it may instead be added inside the argument of the proximal mapping, effectively making the primal update read
$X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1}+\sqrt{2\tau}\xi^{n+1}).$
Since the continuous time limit of both versions is the same SDE, our results on asymptotic behaviour in continuous time in Section 3 hold for either variant. In the discrete time analysis in Section 4, we will focus on the version (1.4) with diffusion as the outer step. Similar results can be proven for the other variant and, where necessary, we will mention how our steps need to be modified to obtain them.